The demand function for x is D(p) = 65 - 2p and the supply function is S(p) = 20 + p. What is the price that should be set to restrict the quantity supplied to 30 units?
Answer #1

S(p) = 20 + p. where S is quantity supplied and P is price of the good.

S = 30

P = S- 20

P = 30 - 20 = 10

At $10, supply = 30 units.